### v2

# DeepInsert: Early Layer Bypass for Efficient and Performant Multimodal Understanding

Abstract
The hyperscaling of data and parameter count in transformer models is yielding diminishing performance improvement, especially when weighed against training costs. Such plateauing underlines a growing need for more efficient finetuning and inference, without sacrificing performance. This is particularly pressing for multimodal learning, where the overhead of processing multimodal tokens alongside language data often limits the practical viability of these systems. In parallel, advances in representation learning and interpretability have deepened our understanding of how such models process and encode information. Notably, recent work has uncovered implicit cross-modal alignment in the deeper layers of large pretrained models. Interestingly, this aligns with our own observations that models naturally defer most cross-modal token interactions to deeper stages of computation. Building on this, we propose a simple modification. Instead of concatenation with the language prompt at the start, we insert multimodal tokens directly into the middle, allowing them to entirely bypass the early layers. Our results with diverse modalities: 1) LLaVA & BLIP for vision, 2) LTU for audio, and 3) MoLCA for molecular data, indicate that our method reduces computational costs during both training and inference, while at the very least, preserving, if not surpassing the performance of existing baselines. Our work has important implications for scaling and composing pretrained models in a resource-efficient manner.
1 Introduction
Advances in large language models (LLMs) [6, 13, 79, 63] have spurred the development of general purpose reasoning models capable of multimodal comprehension in diverse application domains. Although scaling has been crucial for this success [39, 87, 32], the growing awareness of the enormous costs associated with large multimodal language models (MLMs/MLLMs) [74, 30] has placed focus on methods for efficient finetuning of pretrained models and speeding up their inference [33, 43]. For multimodal tasks such as vision-language understanding [67, 44, 50], the use of pretrained unimodal models, rather than training from scratch, has proven effective at reducing training overhead [66]. However, overall efficiency remains a challenge, as multimodal inputs impose additional computational costs during both training and inference, particularly when processed alongside language prompts.
Alongside scaling, a major key to general purpose multimodal training is effective cross-modal alignment. In practice, this is usually achieved with a combination of pretraining stages with paired multimodal data, followed by multitask instruction finetuning for specialized downstream tasks [15]. Preference tuning techniques, such as reinforcement learning from human feedback (RLHF) [64] can produce further improvements [76, 81]. However, most works in this area tend to propose modifications to the training algorithm and/or choice of dataset, while largely treating the pretrained components as replaceable black-box models. To complement this body of work, we begin our investigation with a bottom-up approach by analyzing what happens inside these models. For this qualitative analysis, we rely on pretrained VLMs as a starting point.
Within the literature on analyzing representations and interpretability, our initial inspiration comes from Huh et al. [36], who argue that representations from different unimodal models converge towards a universal platonic representation. Although the platonic hypothesis itself is contentious, the underlying evidence suggests that model representations across modalities show increasing alignment in the deeper layers (a quick glance at Figs. 4, 5 in the Appendix reaffirms it). This, coupled with recalling the inner product mechanism of self-attention, instantly invites the next question: in a multimodal setting, are cross-modal tokens likely to interact more in the deeper layers? Analyzing the attention maps suggests that the bulk of the activity is, indeed, confined to the middle and latter layers (see Fig. 2 and results in concurrent work [38]). All of the preceding exposition thus culminates in our central question: Are the early layers of the LLM really necessary for multimodal token processing, and if not, can we remove redundancy without sacrificing performance? Our key contributions can then be summarized as follows.
  1. 1.
    We show that letting multimodal tokens bypass early layers significantly reduces training and inference compute, with virtually no loss in performance—and in some cases, even marginal gains—in an almost free lunch fashion. We refer to this framework as DeepInsert.
  2. 2.
    We comprehensively demonstrate DeepInsert’s utility across modalities in achieving the desired goals, via replicable experiments on popular open-source MLLMs: vision (LLaVA, BLIP [49, 45]), audio (LTU [28]), and molecules (MolCA [52]). We also reveal a natural performance-efficiency tradeoff (see Fig. 1), enabling practitioners to select models based on their specific needs.
(a) Performance vs runtime in LLaVA (vision)
(b) Performance vs runtime in LTU (audio)
Figure 1: Tradeoff between performance and computational efficiency (inference) in (a) LLaVA v1.5-7B and (b) LTU-7B highlight our two contributions. First, we can reduce multimodal processing (efficiency gains), while either maintaining or improving performance with DeepInsert (see layer 4). Second, it is possible to stay competitive with the baseline, while having significant efficiency gains ($\sim 20$$25\%$ speedup, see layer 12).
The remainder of the paper is organized as follows. Sec. 2 begins by placing our work in the context of broader literature. Sec. 3 motivates our framework, while Sec. 4 details the general notion of the architectural modification, highlighting implementation and design considerations. Experiments in Sec. 5 demonstrate the versatility and effectiveness across a variety of popular open source multimodal models. Sec. 6 concludes with promising directions for broader application. In addition to our more practical contributions, our work hints at novel underlying mechanisms for multimodal processing, which may be of independent interest.
2 Related Work
MLLMs & Alignment: The general architecture of MLLMs considered in our work comprises a pretrained LLM, frozen perceptual encoders for different modalities (image, audio, video, etc. [26, 67, 75, 90]), and a trainable mapping module (can be a lightweight MLP or a larger Perceiver [37]) that aligns the modality with the LLM input space. The LLM prompts consist of text and multimodal tokens, the latter of which are obtained from the encoder via the mapping module. This general regime encompasses a wide range of works in all modalities [1, 45, 88, 3, 49, 28, 25, 60], and thus, our architectural modification is fairly universal in terms of applicability.
Based on where alignment is achieved, our method relates best to the idea of model stitching [42, 61], that is, grafting initial layers of a neural network onto the latter half of another with an affine layer. Our work can thus be interpreted as multimodal model stitching for efficiency, without any additional bells and whistles.
Efficiency & Layer Skipping: Low-rank adapters (LoRA) [34, 16] are commonly used for computationally-efficient finetuning of LLM-based models. The low-rank constraint enables finetuning large models with only a small set of additive parameters, reducing memory usage. Note, however, that inference remains just as costly. An alternative is model compression; an existing model can be effectively distilled, leading to accelerated inference and almost-at-par capabilities [22, 80]. But while these approaches exploit inherent parametric redundancies in LLMs, our approach, as we shall demonstrate, is motivated via functional redundancy and, as such, largely complementary.
The other line of work pertains to layer skipping, which has been employed in LLMs for speculative decoding and speeding up inference [14, 21, 18]. More recently, its application to MLLMs has been considered in [71, 85]. Specifically, Shukor and Cord [71] combine layer skipping with compression and pruning of the base LLMs and demonstrate impressive efficiency gains in their custom set-up. However, it is unclear how well their method scales as their models, when trained in the LLaVA setup, are unable to match baseline performance. On the other hand, concurrent work by Zeng et al. [85] proposes adaptive token skipping to speed up training and inference. However, their token selection is ad hoc and LLaVA-centric. It is unclear if it generalizes to other VLMs or other modalities.
Layer-Skipping Approaches. A number of recent works have proposed layer-skipping or depth-adaptive strategies to improve inference efficiency in large models. Methods such as FlexiDepth [54] and AdaSkip [56] dynamically adjust the number of transformer layers executed based on input complexity, while MoLe-VLA [89] and $\gamma$-MoD [55] introduce mixture-of-expert or modality-specific modules to selectively bypass computation. These approaches, while effective within their respective domains, are often tailored to unimodal LLMs (e.g., FlexiDepth [54], AdaSkip [56]) or require additional architectural modifications for multimodal settings (e.g., MoLe-VLA [89], $\gamma$-MoD [55]). In contrast, DeepInsert does not rely on context-dependent routing or auxiliary modules. Instead, it provides a principled framework for inserting multimodal features at deeper layers, enabling full layer-skipping of the initial stages without degrading performance. Importantly, DeepInsert maintains strong performance across a wide spectrum of modalities—including vision, audio, and molecular data—while preserving the same training setup.
Token-Reduction Approaches. Another line of research focuses on reducing the number of input tokens processed during inference. Representative examples include ST3 [93], Skip-Vision [86], FastV [10], VTW [48], PruMerge [69], MADTP [7], and PACT [17], which prune redundant tokens or learn adaptive token selection strategies. These methods are largely motivated by the redundancy of visual tokens in vision–language models. DeepInsert is conceptually orthogonal to this direction: token-reduction can be applied on top of DeepInsert to further improve efficiency. For instance, we implemented FastV [10] and VTW [48] with DeepInsert and observed that DeepInsert+FastV/VTW maintains performance comparable to or better than baseline+FastV/VTW. This suggests that DeepInsert is particularly robust when combined with token pruning, and can yield additional efficiency gains beyond baseline models. Moreover, training-free methods such as PruMerge [69] also exhibit lower performance declines when applied to DeepInsert models compared to standard LLaVA, reinforcing the complementarity of these two paradigms.
3 Where Multimodal Tokens Matter
This section motivates our proposed framework. As given in our introduction, we first explore the nature of representational alignment in pretrained unimodal models, in the setup of Huh et al. [36]. However, it remains difficult to quantify exactly how this alignment could be beneficial and thus for brevity, we relegate it to Appendix A.1. For concrete evidence, we propose analyzing the attention maps of pretrained MLLMs. The inspiration comes from works on mechanistic interpretability, which suggests that the forward pass of an LLM can be broadly delineated into phases: a) the function determination phase that processes the prompt to determine the task that needs to be executed (for instance, get_capital()) and b) the execution of the function for the given input query (for instance, get_capital(France) = Paris) [31, 57, 59]. These observations lead us to ask whether a similar delineation holds for LLM-based MLLMs: specifically, if we interpret the multimodal tokens as the input query, can we identify where the analogous query execution takes place? Note that query execution would take an inherently different form, since multimodal tokens are generally either prepended, or inserted in the middle of language prompt. Therefore, attention maps are a natural proxy here for understanding how visual tokens move information to the language token for solving multimodal tasks. For this study, we use the LLaVA v1.5 model [49] as our testbed.
Our approach is as follows. We start with a balanced subset of multiple choice type questions (MCQs) drawn from the Visual7W dataset [92], which draws images from COCO [11]. Using MCQ-based questions allows us to first isolate data samples where the model succeeds in answering correctly by comparing the token logit against the choices (A, B, C, or D) and then, obtain the relevant attention activity for these examples, both in just one forward pass. Note that isolating the correct instances is important, since instances of hallucination can significantly alter the attention behavior [38]. The corresponding attention scores from the visual token to the last language token responsible for prediction, as a function of model depth, can then be visualized either by averaging across attention heads, or isolating average behavior for the top-$k$ attention heads per layer. Finally, note that we are interested in layer-relative contribution per visual token. Therefore, we normalize the contribution in each layer by its net contribution across layers. This allows us to cleanly visualize both tokens that contribute high absolute value of attention scores (such as CLIP summary tokens [62]) with those that do not. Finally, for ease of visualization, we randomly subsample image tokens (100 out of 576 tokens) and arrange them left to right in order of increasing positional indices. In the main exposition, we visualize attention with top-$k$ ($k=5$ heads for each layer), in Fig. 2. and relegate the other cases (aggregate over heads, with/without normalization) along with some secondary details, to Appendix A.2.
It is evident from the figure that that vision tokens tend to interact with language predominantly in the intermediate layers. Furthermore, the notion seems to hold for other visualization choices (see Appendix A.2), lending credence to our intuition that vision-to-language information transfer is stronger in intermediate layers and that the earlier layers are somewhat redundant when it comes to performing this effectively. We note that a concurrent line of experimentation reveals similar findings [38], albeit with the motivation to study object hallucination.
Figure 2: Illustration of vision to language token activation in LLaVA-7B. The y-axis represents LLM depth, the x-axis consists of vision tokens arranged in increasing order of positional indices. The score represents the relative layerwise contribution for that token, based on the top-5 attention heads per layer. The figure suggests that the latter layers contain the majority of attention activity.
Remark. In the complete layer-wise depiction (Fig. 6 in the Appendix), there is a strong activation pattern in the first two layers that dominates all other activity. However, we believe that this interaction is unlikely to be the key point of vision-to-language information transfer and we offer two key pieces of evidence for this. Firstly, unlike activation in the middle layers, [4, 38] suggest that this behavior is model-specific (i.e., LLaVA), and is not consistently observed for other choices of VLMs/encoders. Secondly, our own experiments in the Appendix B suggest that even in the base LLaVA model with no pretraining, these layers can be skipped without significant drops in performance (at least to the extent suggested by the activity). Consequently, we do not include the first two layers in Fig. 2 to keep the visualization cleaner.
4 Proposed Framework
Our observations above bring us back to the central question of our study: Do we really need the early LLM layers to parse vision (or more generally, multimodal) information? That is, can we limit multimodal processing to only later layers, without sacrificing performance? To investigate, we propose our simple yet effective modification for general MLLM frameworks: inserting the multimodal tokens directly into the intermediate layers of the LLM. In some sense, one can view it as finding an alternative alignment in the models’ respective semantic representation spaces, akin to model stitching. We call this DeepInsert and illustrate the general approach in Fig. 3. Before jumping into experiments, we first highlight (in our opinion) some important considerations for readers looking to implement this in practice.
(a) Conventional architecture for popular MLLMs.
(b) DeepInsert architecture for MLLMs.
Figure 3: Contrasting our DeepInsert architecture with popular conventional multimodal LLMs. We skip the initial layers for the multimodal tokens to improve computational efficiency while maintaining similar or sometimes better performance.
4.1 Technical Challenges
Although straightforward as an abstraction (Fig. 3), our approach poses several challenges. Recall that while multimodal tokens are inserted in the intermediate layer, language tokens still must be processed by the entire LLM for accurate language processing. Consequently, we must divide the prompt into language and multimodal constituents. The first stage of the forward pass processes just the language tokens. In the second stage, these are recombined within the intermediate layer, before being processed by the latter layers of the LLM. From an implementation point of view, this requires refactoring the LLM’s forward pass as well as the cache structure, so that the model can maintain efficient generation capabilities. Interleaved text and multimodal data raises an additional issue: our proposition of splitting and recombining the prompt may lead to inconsistencies with respect to the positional embeddings in the two stages of the forward pass and thus must be accounted for.
Remark: Note that as an alternative, we can pass the same number of dummy multimodal tokens in the first stage (masked out by a suitable attention mask), before replacing them with the vision tokens. Although this approach eases implementation and may help avoid issues with positional embeddings, it naturally lowers the efficiency gains and, we find, leads to worse performance than skipping.
4.2 Design Considerations
With implementation out of the way, the next question is: in which layer should the multimodal tokens be inserted? If we were to only prioritize efficiency, we would only insert in the last few layers. However, carelessly reducing the number of layers through which the tokens are processed will ultimately lower the capacity of the MLLM to exploit that modality, whether in terms of moving information to language tokens via self-attention or in terms of memory recall within the MLP layers. At present, we sweep this tradeoff by training models with DeepInsert in different layers and comparing their performance. Indeed, characterizing the morphospace and associated Pareto frontier for the tradeoff between efficiency and learning capacity is an intriguing theoretical question.
Nevertheless, woe befall us were we to let our reader depart empty-handed; thus, we offer a path to a compute-friendly heuristic for finding a suitable layer for DeepInsert. It turns out that even without any training, simply loading pretrained MLLM weights onto the DeepInsert architecture and inserting multimodal tokens directly into the middle layers can yield competent MLLMs. And while no longer as competitive, the performance degradations, as one inserts deeper, offer a good proportional reference for training the corresponding DeepInsert model. In other words, if the proxy degradation for a given layer is small enough, the trained DeepInsert will likely match the baseline. We find this to be the case for both LLaVA and LTU models (see Appendix B), and while not perfect, it offers a nice sanity check for the choice of DeepInsert layer, at essentially no extra cost.
5 Experiments
Let us now observe DeepInsert in action. Our experiments are designed to test the framework in a wide variety of settings for: a) different modalities (Vision: LLaVA & BLIP), (Audio: LTU), and (Molecular graph data: MolCA), b) different base LLM sizes (ranging from 350M to 7B parameter models), and c) different LLM architectures (decoder-only & encoder-decoder). To account for variations due to training settings (number and type of GPUs, gradient accumulation, missing data, etc.), we also train the baseline versions for each modality. Experiments on the open-source models i.e. LLaVA, LTU, and MolCA are faithfully reproduced (see Sec. 5.2 for LTU missing data issues) as per the available training code. On the other hand, due to the scale as well as replicability issues with BLIP (unavailable training code, hyperparameters and dataset sampling ratios, takedown of LAION dataset [68, 78]), we compare against the DeepInsert variant in a custom multitask training setup. For brevity, we move this experiment to Appendix C.1.
[…]
Remark. We perform no additional tuning for training the DeepInsert models, proceeding with the same hyperparameters as those used for the baselines. This choice is very deliberate, to emphasize that once we select an insertion layer, DeepInsert variants can easily match baseline performance and yield efficiency gains, while requiring no additional hyperparameter tuning when being adapted to open-source models in the wild.
5.1 Vision
LLaVA (Large Language-and-Vision Assistant) is an end-to-end multimodal model that combines a vision encoder with an LLM, along with a trainable projector MLP that maps visual features onto the LLM embedding space. The model is instruction-tuned in two stages: pretraining for feature alignment and finetuning for multimodal conversational tasks. A key innovation is its use of a ChatGPT-generated long-form instruction-following dataset, enabling LLaVA to excel in tasks requiring deep visual comprehension and reasoning. Given its immense popularity, it represents a natural testbed for evaluating our framework at scale.
We evaluate our method with LLaVA v1.5-7B [49] serving as our baseline, on commonly-used benchmarks for models in that vein. These benchmarks include GQA [35], TextVQA [72], SciQA-IMG-IMG [53], MME [82], POPE [47], MMBench [51], and MM-Vet [83]. For a more detailed discussion of these benchmarks, along with their respective evaluation setups and metrics, the reader may refer to Liu et al. [49]. Additionally, we incorporate two more benchmarks with accuracy as the evaluation metric: AI2D [40] to assess the model’s ability to comprehend and interpret visual information within diagrams, and MMMU [84] to evaluate perception and integration of multimodal information when dealing with subject-specific knowledge and reasoning.
Our results in Table 1 suggest that until layer 4, DeepInsert on average indeed matches the baseline, with the DI-4 matching performance in 2 benchmarks, being marginally outperformed in 4, and outperforming the baseline in the remaining 3 benchmarks. Beyond this point, the efficiency gains come at the cost of performance drops. DI-8 perhaps represents best value, with performance drops within $1\%$ for most datasets, all while skipping $1/4$th of the multimodal compute.
[…]
#Insert Layer GQA TextVQA SciQA-IMG MME POPE MMBench $M$-Vet AI2D MMMU Fwd. pass Finetune
(ms) (hrs:min)
0 (Baseline) 63.1 57.6 69.4 1436.2 86.0 66.1 31.3 56.2 36.1 101.09 45:32
4 (DI) 63.1 57.5 68.7 1483.9 86.0 66.0 32.6 56.0 36.3 93.80 42:01
8 (DI) 62.3 57.1 68.8 1455.5 84.8 65.3 28.8 52.9 34.9 89.08 39:46
12 (DI) 61.1 55.3 65.9 1248.5 83.9 55.2 30.3 52.2 35.6 84.51 36:48
5.2 Audio
LTU (Listen, Think and Understand), introduced by Gong et al. [28], is a popular open-source Audio LLM. The model integrates an Audio Spectrogram Transformer [26], finetuned as an Audio encoder, with a Vicuna-v1.5-7B [12] LLM to enable text-based outputs. A key contributor to its strong performance is the extensive pretraining and finetuning on the OpenAQA-5M dataset, also introduced in the same work. This dataset comprises 1.9 million closed-ended and 3.7 million open-ended audio-question-answer tuples, which helps facilitate the training of models that can both perceive and reason about audio. Note that due to issues with unavailability of the underlying audio (mainly from AudioSet [24]), roughly $10\%$ of the data samples are not usable during training. [^1] This has also been reported in other works such as Ghosh et al. [25]. The training comprises 4 stages: stage-1 trains the audio projection layer with the closed-ended classification and acoustic feature description tasks. In stages 2-4, trainable LoRA adapters are introduced, and the openness of the training task is gradually increased, from only the classification task and acoustic feature description tasks (stage 2), to all closed-ended tasks (stage 3), and finally to all closed-ended and open-ended tasks (stage 4). This perception-to-understanding curriculum eases model transition through the stages, thus helping improve performance.
Consequently, we use LTU as our baseline, and compare DeepInsert on benchmarks utilized by Gong et al. [28], with five audio classification and two audio captioning tasks. Classification accuracies are computed on the eval splits of the ESC50 [65], Vocal Sound [27], and VGG Sound [9] datasets. Similarly, the mean average precision (mAP) classification score is computed on FSD50K [23] and AudioSet [24]. Finally, SPICE scores [2] are used to evaluate the models on captioning tasks using the AudioCaps [41] and Clotho [19] datasets. Please refer to Gong et al. [28] for more details on the evaluation as we use the same process and metrics. We do not find any significant performance differences between the baseline model we trained and the values reported in Gong et al. [28].
Table 2 shows the performance of DeepInsert-LTU as a function of the insertion layer. We observe that insertion at layer $4$ results in the best average performance for both classification and captioning tasks. Somewhat surprisingly, all the models upto DI-24 manage to stay fairly competitive with the baseline, as can be observed by the average performances, hinting at higher redundancies in audio.
Table 2: Performance and runtime of LTU-7B for various audio-language benchmarks. The best performance scores are in bold and the second best are underlined. We clearly see that models trained using multimodal insertion at deeper layer not only match the performance of the baseline, but often exceed it. As expected, the computational time for both training and inference drops as we insert deeper into the network (training was done on $\timesH2$00 141 GB GPUs).
#Insert layer ESC50 VS VGG FSD AudioSet Classif. AudioCaps Clotho Cap. Fwd. pass Train time
(Acc) (Acc) (Acc) (mAP) (mAP) Avg. (SPICE) (SPICE) Avg. (ms) (hrs:min)
0 (Baseline) 85.45 56.42 49.64 46.51 19.13 51.43 16.55 11.77 14.16 8.20 17:30
4 (DI) 86.10 57.98 49.80 44.50 18.94 51.46 16.46 12.06 14.26 7.80 16:16
8 (DI) 84.45 57.34 48.87 46.57 18.97 51.24 16.32 12.01 14.17 7.37 15:38
12 (DI) 85.10 56.42 49.95 44.73 19.23 51.09 16.37 11.82 14.10 6.93 15:04
16 (DI) 85.90 52.74 49.64 42.93 19.15 50.07 16.62 11.71 14.17 6.49 14:26
24 (DI) 82.30 53.41 50.39 44.92 19.51 50.11 16.30 11.88 14.09 5.63 13:22
5.3 Molecular Data
MolCA (Molecular Graph-Language Modeling with Cross-Modal Projector and Uni-Modal Adapter) enables language models to process both molecular graphs and text by bridging the two modalities [52]. The training pipeline is almost identical to InstructBLIP: it uses a Cross-Modal Projector (a QFormer with a SciBERT initialization [5]) to translate information from molecular graph features (obtained from a graph encoder) into learnable query tokens that are fed into the LLM. Additionally, a unimodal low-rank adapter is used to fine-tune the LLM efficiently. The primary tasks considered include molecular-captioning and IUPAC name prediction. To better suit molecular data, Galactica model variants (125M and 1B) are considered as the pretrained LLMs [77].
For our experiments, we follow the exact same training regime as Liu et al. [52]; the pretraining stage 2 for aligning 2D molecular graphs to texts via language modeling (10 epochs on their curated Pubchem324k pretrain split [52]), followed by finetuning unimodal adapters for downstream tasks. Since stage 1 does not involve the LLM, we simply start with the Stage 1-pretrained QFormer for both models. We focus on the primary task of the paper, that is, the molecular captioning experiments, using Pubchem324k and CheBI-20 [20]. For the LLM, we rely on the larger Galactica-1B [77] variant, with 24 LLM layers. The LoRA configurations for finetuning on PubChem324k and CheBI-20 datasets follow the same configuration as in Liu et al. [52]. The DeepInsert modification injects the QFormer in different layers and we report their performance based in the same captioning evaluation metrics, as used in Liu et al. [52]. For both datasets, we train the model for the reported 100 epochs and check scores every 10 epochs. Since final performance does not correspond with best performance for the baseline, we report the best model performance obtained during the training run.
Our results in Table 3 are consistent with our findings in the other modalities; DI-9 and DI-12 can match and/or outperform the baseline model with ease, on both datasets. Due to the inherent drawbacks of captioning metrics used in the paper, it is harder to quantify the marginal gains as such. Notwithstanding, the result does indicate the redundancy of multimodal processing also for molecular data, up to even $50\%$ (DI-12 ends up using only uses half the LLM layers for the multimodal tokens).
Table 3: Molecular captioning performances on PubChem324k and CheBI-20 datasets. Bold and underline indicate best and second-best performance, respectively. Both DeepInsert models (DI-9, DI-12), achieve similar or better performance than the baseline, using roughly only half the the LLM for multimodal processing.
#Insert Layer BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
0 (Baseline) 58.3 49.5 66.1 51.4 59.9 61.3
9 (DI) 58.6 49.8 66.4 51.7 60.3 61.6
12 (DI) 58.2 49.5 66.2 51.6 60.1 61.3
(a) Molecular Captioning on the CheBI-20 dataset for Galactica-1.3B based MolCA models.
#Insert Layer BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
0 (Baseline) 34.1 26.4 48.5 33.9 43.1 41.8
9 (DI) 34.8 27.2 49.4 34.7 43.8 43.0
12 (DI) 35.3 27.7 49.7 35.3 44.3 43.4
(b) Molecular Captioning on the Pubchem324k dataset for Galactica-1.3B based MolCA models.
5.4 Reflections on Performance vs. Efficiency
Our results validate the general efficacy of DeepInsert in reducing multimodal redundancy. While the efficiency improvements naturally follow from the specific choice of insertion layer, we include concrete evidence in terms of training and inference run-times for some of the larger models in Tables 1 and 2. This trade-off can be better visualized as a function of insertion layer in Fig. 1, where we train models for LLaVA and LTU with insertions upto layer 24, report their respective performances on a representative set of benchmarks and overlay it against the average inference forward pass time. Furthermore, the interested reader can find a more careful breakdown of gains in terms of expected number of FLOPs, and how it lines up against what is observed in Appendix D.
The actual numbers aside, perhaps the more pertinent question here is why certain models and modalities seem to respond better to DeepInsert than others. Specifically, while both LTU and MolCA can be reduced to using only half the LLM with virtually no degradations, LLaVA performance seems to degrade faster after layer 8. Here we speculate why it may be the case. Firstly, we note that one contributing factor could be the number of multimodal tokens. In general usage with some 100 textual prompt tokens, while both LTU and MolCA rely on 32 multimodal tokens, LLaVA v1.5 utilizes all 576 vision tokens from the vision encoder. Consequently, we expect this to lower the redundancy in layers due to the proportional increase in multimodal load. This is likely also exacerbated by the respective training exposures (the recipes for training MolCA and LTU require significantly higher ratios of data and/or training iterations for multimodal alignment, when compared against the number of multimodal tokens). Note that the there is significant scope to reduce the number of vision tokens, as noted in works that propose more token-efficient versions of LLaVA [70, 46]. Therefore, repeating our DeepInsert experiments with those variants could yield much better proportional gains than the baseline.
6 Discussion
In this work, drawing inspiration from alignment and interpretability literatures, we introduced DeepInsert. We then back it up with concrete experiments for different MLLMs to show that our framework can produce (a) efficiency gains with no effects on performance up to an extent, and (b) follow it up with a reasonable performance-efficiency trade-off. Perhaps the greatest strength of DeepInsert lies in its versatility, in that, it can be successfully applied to off-the-shelf MLLMs without any additional tuning. And while layer selection remains MLLM-specific, the heuristic provided in Appendix B can help guide that choice, before any training.
Our findings open up several lines of research. The next step in advancing our work is to explore its empirical limits, with a focus on scalability across model size and dataset diversity. A natural extension would be to apply it to video MLLMs, where long-form multimodal content presents substantial computational challenges. Additionally, a more rigorous characterization of why certain modalities and frameworks offer different performance-efficiency trade-offs could serve as a two-fold key for improving both our (still nascent) mechanistic understanding of multimodal processing in transformers as well as the practical suite of methods for combining pretrained models into larger systems. Finally, on the theoretical front, if our intuition on treating multimodal tokens as a query (in Sec. 3) can be firmly established, it can perhaps help isolate how multimodal information is utilized, thus building towards mechanisms behind multimodal factual retrieval, akin to language counterparts [31, 57, 59].
7 Limitations & Societal Impact
We only tested our approach on a selective subset of MLLMs. Secondly, despite the offered heuristic, we are unable to offer concrete guarantees on performance without retraining the model, which does expend compute. Finally, since we stick to the baseline evaluation setups, we do inherit their natural deficiencies (for instance, the known drawbacks of captioning performance metrics in MolCA).
In terms of societal impact, since our work focuses on (pretrained) MLLMs, we suffer from the standard pitfalls associated with such models, including hallucination, equitable access, lack of accountability, and data attribution. At the same time, our proposed method enhances efficiency, potentially lowering computational costs and energy requirements for multimodal model development while maintaining performance, leading to positive societal, economic, and environmental impacts.
Appendix A Additional Experiments: Interpretability
A.1 Vision-Language Alignment in Pretained Unimodal Models
We begin with some formal exposition to characterize the notion of alignment, borrowed from Huh et al. [36]. A representation is a function $\mathbf{f}:\mathcal{X}\to\mathbb{R}^{n}$ that assigns a feature vector to each input in some data domain $\mathcal{X}$. A kernel, $\mathbf{K}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, characterizes how its corresponding representation measures distance/similarity. It is defined as $\mathbf{K}(x_{i},x_{j})=\langle\mathbf{f}(x_{i}),\mathbf{f}(x_{j})\rangle,$ where $\langle\cdot,\cdot\rangle$ denotes the inner product, $x_{i},x_{j}\in\mathcal{X}$, and $\mathbf{K}\in\mathcal{K}$. Finally, a kernel-alignment metric, $\mathbf{m}:\mathcal{K}\times\mathcal{K}\to\mathbb{R}$, measures the similarity between two kernels, that is, how similar is the distance measure induced by one representation to the distance measure induced by another. As in Huh et al. [36], we rely on a mutual nearest-neighbor metric that measures the mean intersection of the $k$-nearest neighbor sets induced by two kernels (normalized by $k$).
In essence, we measure the distance between the kernels induced by the representation spaces of language and vision. First, these representations are obtained by passing paired image-caption data through the respective pretrained models. Then, we rely on a mutual $k$-nearest neighbor alignment metric to estimate the alignment between the corresponding feature spaces. We want to understand how alignment varies as a function of model depth, with the intuition that layers capturing “world semantics" are likely to align better, even across modalities. For the LLaVA v1.5-7b model, its starting components include the vision encoder, clip-vit-large-patch14-336 [67], and the decoder-only LLM vicuna-v1.5-7b model [91]. To compute the alignment metric, we use $k=10$ nearest neighbors over 1024 samples from WIT (Wikipedia-based Image Text; [73]); additional details can be found in Huh et al. [36].
The alignment plot in Fig. 4 is in line with our expectations: the penultimate layer of the vision model aligns more with the latter layers of the LLM. This may point to some shared semantics in the deeper layers of the respective models, and can perhaps be exploited in designing alignment frameworks for VLMs. Note that this alignment is more general. Fig. 5 shows a similar trend in DINO-based encoders [8]. This is remarkable as it points to a more general preexisting alignment, since DINO vision encoders are—unlike CLIP—trained in a self-supervised language-agnostic fashion.
Figure 4: Visualization of intrinsic alignment between pretrained vision (CLIP-based ViT) and language models (Vicuna v1.5), as a function of depth. The vertical axis represents the LLM layer indices, whereas the horizontal axis does the same for the vision encoder. The key point is that for the representations from the latter vision layers (which are generally used as embeddings), alignment becomes stronger as one moves deeper into the LLM.
Figure 5: Visualization of intrinsic alignment between pretrained vision (DINO-based ViT) and language models (Vicuna v1.5), as a function of depth. The trend is similar to that for CLIP-based models, in for the representations from the latter vision layers (generally used as embeddings), alignment becomes stronger as one moves deeper into the LLM.
A.2 Where Visual Tokens Matter
Here, we present the complete list of visualizations, to comprehensively demonstrate the middle layer activation. We begin by visualizing the attention activation for all layers. As we can see, the activation in the middle, while still prominent, is overshadowed by the flurry in the first two layers. Nevertheless, as mentioned in the main paper, this is likely of little import for analyzing the general trend. We also demonstrate that the visualization when done over the average behavior of activation heads as opposed to top-5 (see Fig. 7), or without normalization (see Fig. 8) does not make a difference in terms of the nature of our conclusions. For the latter, even though we see the high magnitude CLIP tokens dominate the attention matrix, their activity is also more middle layers heavy.
Figure 6: Full illustration of vision to language token activations (including the first two layers), in the forward pass of a LLaVA-v1.5-7B model.
Figure 7: Illustration of attention activity, averaged over all heads.
Figure 8: Illustration of attention activity, without token specific normalization.
Appendix B Layer Selection Heuristic
The key idea is as follows: if there is intrinsic evidence that multimodal tokens are not being used in the early layers, can we skip them even in baseline pretrained models? Our somewhat surprising find is that it is possible to do so, without any finetuning. Specifically, we load the pretrained model weights of the MLLM (LLaVA or LTU) onto the DeepInsert variant, by simply letting the multimodal adapter insert tokens into the latter layer.
Skip Layers (LLaVA) POPE
0 86.0
DI-4 84.4
DI-8 77.9
DI-12 56.3
DI-16 47.8
Table 4: Performance across layers, for baseline LLaVA pretrained weights, on POPE benchmark.
Skip Layers (LTU) ESC50 VS
0 81.30 56.39
DI-4 81.70 56.36
DI-8 74.50 41.21
DI-12 66.40 31.05
DI-16 32.35 32.36
Table 5: Performance across layers, for baseline LTU pretrained weights, ESC50 and VS audio benchmarks.
From Table 4 and 5, we note that both models retain competency even when layers are skipped. While the performance drops in LLaVA are modest yet noticeable, LTU surprisingly even improves for DI-4, suggesting a strong implicit redundancy, at least for classification benchmarks. We also note that these trends are also indicative of which model/modality are likely to be more susceptible to DeepInsert.
We did however notice that with both models, the lengths of text generation seems to monotonically decrease (to the point that the evaluation benchmark for captioning seems to fail for LTU). Therefore, direct insertion without retraining degrades performance on generative benchmarks and there is still a need to retrain models from scratch.
Appendix C Additional Experiments: MLLMs
C.1 BLIP
BLIP (Bootstrapped Language-Image Pretraining) is a pioneering approach for vision-language modeling [44]. It introduces a lightweight Querying Transformer (QFormer) that extracts information from the embeddings of the vision encoder. Vision information is transferred into a set of learnable query tokens (32 embeddings of dimension 768) via the QFormer’s cross-attention layers, which are then fed into the LLM along with the language prompt. Further improvements by: a) scaling the LLM and pretraining dataset [45], and b) multi-task instruction finetuning [15] allows the model to solve a variety of tasks like image captioning, visual question answering, and multimodal reasoning.
In our experiment, we will consider one such model to demonstrate the potential of DeepInsert. Specifically, we consider the VLM’s construction with FlanT5-base [13] (encoder-decoder architecture, 350M parameter model) as our LLM and the Eva-clip-g/14 [75], a CLIP-based vision transformer, for our frozen image encoder. The QFormer follows the default architecture and initialization as in Dai et al. [15]. Since FlanT5-base has an encoder-decoder architecture, insertion bypasses the encoder entirely. Instead of feeding the Qformer outputs (query tokens) as inputs to the encoder, we feed them directly into the decoder by prepending them to the language tokens. The choice of bypassing the encoder is intuitive, since the encoder output space is likely to capture the semantics of the language prompt and therefore align better with the image semantics captured in the QFormer tokens.
Note that we do not intend to follow the training recipe for the InstructBLIP models. Indeed, it is impossible to replicate the model for two reasons: a) lack of information on training recipe, including learning rate schedules and dataset sampling ratios, and b) parts of the LAION dataset, the key component for pretraining, was taken offline so as to comply with U.S. federal law, cf. Thiel [78]. Furthermore, the high-level recipe calls for multi-epoch pretraining on a dataset with roughly 130M image-text pairs, which are beyond the scope of this work. Instead, we define a smaller setup, which still allows us fair testing of DeepInsert but with a reasonable compute budget. Recall that the end-to-end training phase for InstructBLIP involves two steps: the pretraining step with image-caption pairs, followed by the multi-task instruction finetuning step. Consequently, we structure the experiment in a similar vein, by first pretraining efficiency, followed by a comparison post multi-task instruction finetuning.
C.1.1 Pretraining Efficiency
Reflecting the InstructBLIP setup, we consider image-caption pretraining with the COCO captions dataset [11] as the first stage. The goal of this experiment is to compare the level of alignment between the QFormer and the LLM, as training proceeds. Specifically, we measure the best possible captioning score achieved within a fixed number of epochs (1, 10 and 20) for each framework. This is done in accordance with different hyperparameter and learning rate configurations available in the BLIP repository and for each variant, we report the highest performance across configurations. As shown in Fig. 9, we find that our framework not only achieves better peak performance, but also does it in much fewer epochs. We limit our report to 20 epochs as both models achieve over 95% of their respective peak scores within that frame (the peak is achieved at around 50 epochs).
Figure 9: Pretraining comparison for default (QFormer) vs. grounded (Ours) pipelines. Higher score indicates that DeepInsert achieves better alignment, and in much fewer epochs.
C.1.2 Multi-task Training
Next, we consider instruction finetuning for two tasks: Captioning and Visual Question Answering (VQA). For each task, we sample from a fixed set of text prompts in every training iteration.
To emulate InstructBLIP’s training stages, we first pretrain with image-caption pairs for 20 epochs, and then perform multi-task instruction finetuning for captioning and VQA for 15 combined epochs (uniformly sampling from the datasets without replacement). For captioning, we consider the COCO captions dataset as before and for visual question answering, we consider VQAv2 [29] for training. Instruction finetuned VLMs can demontrate remarkable zero-shot abilities on unseen datasets, and so we also consider OKVQA (Outside Knowledge VQA) [58] for zero-shot evaluation. Coco Captions and VQAv2 share the same train-val-test image split that aims to prevent train-test information leakage for a fair evaluation.
Results are given in Table 6. We find that for the designated tasks, our method achieves significant improvements over the standard QFormer. We also see a remarkable gap in zero-shot OKVQA performance, with the grounded model even outperforming the reported performance of some of the largest BLIP-2 models [45]. (InstructBLIP includes OKVQA during training, so no direct comparison is possible.)
Model COCO Captions (BLEU-4) VQAv2 (%) OKVQA (0-shot) (%)
Default (QFormer) 20.9 55.4 28.8
Grounded (Ours) $\bm{36.2}$ $\bm{66.8}$ $\bm{39.0}$
Table 6: Performance comparison for BLIP-inspired multi-task instruction finetuned models, including 0-shot evaluation on unseen dataset. We find that in our custom setup, bypassing the encoder and inserting directly to the decoder can yield significant performance gains.
C.2 Training Loss Comparison on LLaVA
Figure 10: Comparison of training losses of LLaVA finetuning stage, for the baseline model against best performing grounded model (layer-4), revealing a tendency to achieve lower next token prediction loss and potentially better learning.
Appendix D Efficiency Analysis
The number of FLOPs performed by a transformer model in a forward pass can be analytically expressed as a function of the transformer’s various hyperparameters such as model hidden dimension, feedforward layer dimension, and number of layers. In this section, we extend the analysis to DeepInsert and compare the theroetically expected FLOPs with the empirically measured runtime. Let the transformer have $N$ layers with a hidden dimension of $d_{\text{model}}$ and a feedforward dimension of $d_{\text{ff}}$. Let the number of heads for multihead attention be given by $n_{\text{head}}$. Assume the number of multimodal tokens is $L_{\text{mm}}$ and the number of text tokens in a certain forward pass is $L_{\text{text}}$. Let the DeepInsert layer be given by $N_{\text{DI}}$.
The total number of FLOPs (additions plus multiplications) in one forward pass can be divided into the following parts: (i) the projections into Query ($Q$), Key ($K$), and Value ($V$) matrices, (ii) the self attention mechanism, (iii) the feedforward layers, and (iv) other computations such as activations, layer norm, etc. Since (iv) is negligible compared to (i), (ii), and (iii), we ignore it here.
Expanding on (i), the input $X\in\mathbb{R}^{L\times d_{\text{model}}}$ needs to be multiplied with projection matrices $W_{q},W_{k},W_{v}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ to give $Q,K,V$ respectively. For an input sequence of length $L$, each $XW$ matrix multiplies require $2Ld_{\text{model}}^{2}$ operations including both additions and multiplications, giving us
$\text{Projection FLOPs}=6Ld_{\text{model}}^{2}.$ (1)
Moving to (ii), we first need to compute the $QK^{\top}$ matrix, and since $Q$ and $K$ are both in $\mathbb{R}^{L\times d_{\text{model}}}$, this requires $n_{\text{head}}\times 2L^{2}\frac{d_{\text{model}}}{n_{\text{head}}}=2L^{2}d_{\text{model}}$ FLOPs. We then apply a row-wise softmax to get the $L\times L$ attention matrix for each head, which requires another $2L^{2}n_{\text{head}}$ FLOPs. Next, the $L\times L$ attention matrices are multiplied with the value matrices which takes up another $2L^{2}d_{\text{model}}$ FLOPs across all heads. Finally, the output projection takes up another $2Ld_{\text{model}}^{2}$ FLOPs. Adding all these up, and using the approximation $n_{\text{head}}<<d_{\text{model}}$, we have
$\text{Attention mechanism FLOPs}=4L^{2}d_{\text{model}}+2Ld_{\text{model}}^{2}.$ (2)
Computing (iii) is more straightforward as the feedforward layers are pairs of linear layers, with dimensions $d_{\text{model}}\times d_{\text{ff}}$ and $d_{\text{ff}}\times d_{\text{model}}$. This gives us
$\text{Feed forward FLOPs}=4Ld_{\text{model}}d_{\text{ff}}.$ (3)
Adding (1), (2), and (3) gives us our final analytical expression for the number of FLOPs per layer when the sequence length is $L$:
$\text{FLOPs per layer for $L$ tokens}=8Ld_{\text{model}}^{2}+4L^{2}d_{\text{model}}+4Ld_{\text{model}}d_{\text{ff}}.$ (4)
Using DeepInsert, the first $N_{\text{DI}}$ layers only see $L_{\text{text}}$ tokens, while the remaining $N-N_{\text{DI}}$ layers see all the $L_{\text{mm}}+L_{\text{text}}$ tokens. This give us the total number of FLOPs for deep insertion at layer $N_{\text{DI}}$:
$\text{FLOPs}(N_{\text{DI}})=N$ $\Big(8(L_{\text{text}}+L_{\text{mm}})d_{\text{model}}^{2}+4(L_{\text{text}}+L_{\text{mm}})^{2}d_{\text{model}}+4(L_{\text{text}}+L_{\text{mm}})d_{\text{model}}d_{\text{ff}}\Big)$
$-N_{\text{DI}}\Big(8L_{\text{mm}}d_{\text{model}}^{2}+4(2L_{\text{text}}+L_{\text{mm}})L_{\text{mm}}d_{\text{model}}+4L_{\text{mm}}d_{\text{model}}d_{\text{ff}}\Big).$ (5)
As we see from (5), the number of FLOPs decreases monotonically as a function of $N_{\text{DI}}$. Further, a larger number of multimodal tokens $L_{\text{mm}}$ leads to a steeper decrease in FLOPs as we insert deeper into the network.
To validate (5), we empirically measure the runtime of LLaVA on 80 GB A100 GPUs and plot the both the runtime and the theoretically estimated FLOPs against the insertion layer in Fig. 11. As we see in Fig. 11, there is a very close match between the theoretically estimated FLOPs of (5) and the empirically measured runtime of the forward pass of the model, suggesting that (5) can be a useful proxy for the computational efficiency of the model when insertion is done at various deep layers.
Figure 11: Comparing the model runtime with the theoretically estimated FLOPs from (5). As we see, both curves match very closely.